OpenStudy (anonymous):

Can anybody help with this word problem please? Sheila earns a basic wage of $8 per hour. Under certain conditions, she is paid $12 per hour. The most that she can earn in one week is $400. a. Write an inequality that describes her total weekly wages for x hours at $8 per hour and for y hours at $12 per hour. b. Graph the inequality. c. What is the maximum number of hours that Shelia can work for $8 per hour? for $12 per hour? Thank you! =)

OpenStudy (anonymous):

x+y<= 400

OpenStudy (anonymous):

a. x+y less than or equal to 400 b. 8/hr = 50 hours 12/hr = 33 hours

OpenStudy (anonymous):

Thank you @serenity3 and @julian25!